# More precise edge detections

### v1

@@ Abstract
Image Edge detection (ED) is a base task in computer vision. While the performance of the ED algorithm has been improved greatly by introducing CNN-based models, current models still suffer from unsatisfactory precision rates especially when only a low error toleration distance is allowed. Therefore, model architecture for more precise predictions still needs an investigation. On the other hand, the unavoidable noise training data provided by humans would lead to unsatisfactory model predictions even when inputs are edge maps themselves, which also needs improvement. In this paper, more precise ED models are presented with cascaded skipping density blocks (CSDB). Our models obtain state-of-the-art(SOTA) predictions in several datasets, especially in average precision rate (AP), which is confirmed by extensive experiments. Moreover, our models do not include down-sample operations, demonstrating those widely believed operations are not necessary. Also, a novel modification on data augmentation for training is employed, which allows noiseless data to be employed in model training and thus improves the performance of models predicting on edge maps themselves.
@@ III-A The SDPED model
Figure 2:  The SDPED model: Top path is the main path of the model. Typically, the first 2 convention layers output 32 and 64 features, while the number of input and output features in the last two convention layers are all 64. $n+2$ layers, including the feature extractor block, each CSDB, and the last block, would provide middle results, where each would be fused into 21 features. All middle results are concatenated to obtain $21\times(n+2)$ features for the final fusing layer (the button path). Button path is the fusing block, where the $21\times(n+2)$ features from the main path goes through one $3\times 3$ convention block and then two $1\times 1$ convention blocks. The numbers of the output features of such layers are 256, 512, and 1, respectively.
@@ V Discussion
On the other hand, our model could be more versatile compared with previous ones. As pointed out in [41], previous models obtained better scores on some datasets usually with the sacrifice on the scores on others. For example, the experiments in the above section show that HED and RCF can outperform others on BRIND but are outperformed by others on UDED and BIPED2, and BDCN could perform best on UDED but not on other datasets, while Dexi and RIND provide best results on MDBD and BIPED2 but not on the other two. However, our model could obtain the best results on all the three benchmarks on most datasets, namely BRIND, UDED, and MDBD, while significantly outperforming previous models on AP with similar scores on other benchmarks, on BIPED2. Therefore, our model could be more flexible on multiple datasets.
As for obtaining noiseless data for ED tasks, it could be a long-standing problem since edges do not have a formal definition. Although works have tried to define it, all of them fail to distinguish edges and textures, as far as we know [^13] In another view, if edges can be defined formally, traditional algorithms could have solved ED problems.. Therefore, datasets for ED tasks are highly dependent on human labels, which are unavoidable noise. Although Section III B provides the first trick to obtain noiseless data, it is worth investigating further how to obtain such data.

### v3

@@ Abstract
Image Edge detection (ED) is a base task in computer vision. While the performance of the ED algorithm has been improved greatly by introducing CNN-based models, current models still suffer from unsatisfactory precision rates especially when only a low error toleration distance is allowed. Therefore, model architecture for more precise predictions still needs an investigation. On the other hand, the unavoidable noise training data provided by humans would lead to unsatisfactory model predictions even when inputs are edge maps themselves, which also needs a solution. In this paper, more precise ED models are presented with cascaded skipping density blocks (CSDB). Our models obtain state-of-the-art(SOTA) predictions in several datasets, especially in average precision rate (AP), over a high-standard benchmark, which is confirmed by extensive experiments. Also, a novel modification on data augmentation for training is employed, which allows noiseless data to be employed in model training for the first time, and thus further improves the model performance. The relative Python codes can be found on https://github.com/Hao-B-Shu/SDPED.
@@ III-A The SDPED model
Figure 2:  The SDPED model: Top path is the main path of the model. Typically, the first 2 convention layers output 32 and 64 features, while the number of input and output features in the last two convention layers are all 64. $n+2$ layers, including the feature extractor block, each CSDB, and the last block, would provide middle results, where each would be fused into 21 features. All middle results are concatenated to obtain $21\times(n+2)$ features for the final fusing layer (the bottom path). Bottom path is the fusing block, where the $21\times(n+2)$ features from the main path goes through one $3\times 3$ convention block and then two $1\times 1$ convention blocks. The numbers of the output features of such layers are 256, 512, and 1, respectively.
@@ V-B How big SDPED is?
V-C Do it effective in varied ED datasets?
On the other hand, our model could be more versatile compared with previous ones. As pointed out in [41], previous models obtained better scores on some datasets usually with the sacrifice on the scores on others. For example, the experiments in the above section show that HED and RCF can outperform others on BRIND but are outperformed by others on UDED and BIPED2, and BDCN could perform best on UDED but not on other datasets, while Dexi and RIND provide best results on MDBD and BIPED2 but not on the other two. However, our model could obtain the best results on all the three benchmarks on most datasets, namely BRIND, UDED, and MDBD, while significantly outperforming previous models on AP with similar scores on other benchmarks on BIPED2. Therefore, our model could be more flexible on multiple datasets.
V-D Limitation and Future works
Despite the effectiveness of SDPED models with fewer coefficients, their inference speed and GPU consumption might need improvement. Since SDPED models are implemented on the full image resolution, the memory cost could be high and the inference speed might be lower than previous models. In this paper, we mainly focus on improving the performance of ED models over a high-standard benchmark but pay less attention to accelerating them. But in practice, it would be a trade-off between accuracy and consumption. The acceleration schemes and the reduction of the memory are left to the future works.
On the other hand, obtaining noiseless data for ED tasks could be a long-standing problem since edges do not have a formal definition [^13] In another view, if edges can be defined formally, traditional algorithms could have solved ED problems.. Therefore, datasets for ED tasks are highly dependent on human labels, which are unavoidable noise. However, most problems in the model prediction might be traced back to data, which was also indicated in the experiments in Section IV, where most models perform better on UDED dataset with the highest quality even if the dataset contains the least number of training images. Although Section III $B$ provides the first trick to obtain noiseless data, it is worth investigating further how to obtain more datasets with less noise.